You can bike 14 miles per hour. How far can you bike In 2 1/2 minutes?
Answer by stanbon(75887)

You can bike 14 miles per hour.
How far can you bike In 2 1/2 minutes?
-----
2 1/2 minutes is (5/2)/60 = 1/24 hr.
----
Your Problem:
distance = rate*time
distance = 14*(1/24) = 7/12 mile
===================================
cheers,
Stan H.
===========